In a simplistic model of the hydrogen atom, the electron orbits the proton in a circle of radius 53 pm. What is the orbital period of the electron, in seconds, if the force responsible for the proton-electron attraction is electric?
Answer #1


m = mass of electron


v = speed of electron = 2*pi*r/T


e = charge = 1.6*10^-19

r = radius


electric force Fe = k*e^2/r^2

centripetal force Fc = m*v^2/r

Fe = Fc

m*v^2/r = k*e^2/r^2

m*v^2 = k*e^2/r


m*4*pi^2*r^2/T^2 = k*e^2/r

9.11*10^-31*4*pi^2*(53*10^-12)^2/T^2 = (9*10^9*(1.6*10^-19)^2)/(53*10^-12)


T = 1.52*10^-16 s